### v1

# Geometric quantum thermodynamics: A fibre bundle approach

Abstract
Classical thermodynamics is a theory based on coarse-graining, meaning that the thermodynamic variables arise from discarding information related to the microscopic features of the system at hand. In quantum mechanics, however, where one has a high degree of control over microscopic systems, information theory plays an important role in describing the thermal properties of quantum systems. Recently, a new approach has been proposed in the form of a quantum thermodynamic gauge theory, where the notion of redundant information arises from a group of physically motivated gauge transformations called the thermodynamic group. In this work, we explore the geometrical structure of quantum thermodynamics. Particularly, we do so by explicitly constructing the relevant principal fibre bundle. We then show that there are two distinct (albeit related) geometric structures associated with the gauge theory of quantum thermodynamics. In this way, we express thermodynamics in the same mathematical (geometric) language as the fundamental theories of physics. Finally, we discuss how the geometric and topological properties of these structures may help explain fundamental properties of thermodynamics.
I Introduction
The universal character of thermodynamics stems from the type of physical questions it is designed to address. In the thermodynamic limit, experimentally accessible observables probe only coarse-grained averages of the underlying microscopic dynamics, since atomic-scale details lie beyond direct measurement. Such coarse-graining gives rise to the thermodynamic state variables and results in a robust theoretical framework that constrains macroscopic processes largely independently of microscopic details [1]. An essential consequence of this limit is the effective suppression of fluctuations. For small-scale or quantum systems, by contrast, fluctuations play a central role, and under far-from-equilibrium driving the standard thermodynamic description breaks down. In such regimes, one must employ alternative approaches, such as stochastic thermodynamics [2] and information theory [3].
Despite significant advances since Alicki’s pioneering work [4], quantum thermodynamics remains incomplete [5]. Foundational questions persist regarding the definitions of work, heat, and entropy in the quantum regime; see [6, 2] for recent discussions.
In contrast to thermodynamics, which emerges from coarse-grained descriptions of high-dimensional systems, modern fundamental physics—the standard model and general relativity—is formulated in terms of gauge theories [7, 8, 9]. A gauge theory is defined by a Lagrangian that is invariant under local transformations acting on internal degrees of freedom; the requirement of local invariance demands the introduction of gauge fields that mediate interactions.
Geometrically, gauge theories are expressed in the language of fibre bundles: gauge fields correspond to connections on principal bundles, and their field strengths correspond to the associated curvatures. The base manifold represents spacetime, and the fibres encode the internal symmetry group [7, 8, 9, 10, 11, 8]. Although gauge theories originated in particle physics, they appear in diverse contexts, including low Reynolds number hydrodynamics [12, 13], condensed matter physics [14], quantum information [15], and finance [16], to mention just a few.
Classical thermodynamics has also been explored from this gauge-theoretic perspective. Treating the thermodynamic force as a gauge field, Ref. [17] analysed the Onsager relations within nonequilibrium thermodynamics, while Ref. [18] revealed a connection between $U(1)$ lattice gauge theories and nonequilibrium processes. Moreover, by examining the symmetry under local rescaling of probabilities in Markovian dynamics on finite graphs, Ref. [19] proposed interpreting thermodynamic forces as gauge potentials.
Based on information theory, Refs. [20, 21] cast quantum thermodynamics as a gauge theory of the thermodynamic gauge group. Classical thermodynamics relies on coarse-graining, whereas quantum mechanics assumes full control over microscopic degrees of freedom, yielding an excess of information from a thermodynamic standpoint. The thermodynamic gauge group removes this redundant information, playing a role analogous to the elimination of microscopically irrelevant details in classical thermodynamics. Within this formalism, thermodynamic quantities arise uniquely from gauge invariance, which is one of the most fundamental principles of modern physics.
The gauge structure introduced in [20, 21] is particularly rich, featuring a time-dependent gauge group whose geometric properties remain unexplored. The goal of this work is to clarify the underlying geometry of this theory and to construct the associated fibre-bundle structures in a mathematically rigorous way. This puts thermodynamics on the same mathematical structure as the fundamental physical theories. This may shed new light on the foundations of classical and quantum thermodynamics.
The paper is structured as follows. In Sec. II, we review the gauge theory of the thermodynamics group developed in Refs. [20, 21]. Section III presents how gauge theories can be formulated in terms of the geometry of principal fibre bundles and introduces some common examples from field theory to pave the way for the construction of our theory, which is done in Sec. IV. Finally, in Section V, we discuss how our geometric theory may help in understanding thermodynamics and provide new mathematical insights into the theory by using geometrical and topological techniques, as well as proposing future directions for investigation. Additionally, in Appendix A, we revise the mathematics of fibre bundles, principal bundles, and Lie theory, which are used extensively throughout this paper, in order to make our discussion more self-contained.
II Quantum thermodynamics as a gauge theory
Classical thermodynamics is a theory concerned with the macroscopic bulk properties of large systems, where fluctuations vanish. The theory focuses on the average behaviour of the microscopic degrees of freedom (justified by the central limit theorem), while ignoring the microscopic configurations. This means that, in classical thermodynamics, we have limited access to information, which is referred to as coarse graining [1].
In trying to adapt useful concepts of thermodynamics to the quantum realm, quantum thermodynamics has recently arisen as a powerful theoretical framework [6, 2]. In great contrast to its classical counterpart, however, quantum thermodynamics is usually based on a significant amount of information about the system being studied. For instance, thermodynamic protocols are frequently considered in which knowledge of the density operator is mandatory. This is why quantum thermodynamic quantities are often defined in terms of informational measures [6, 3].
In an attempt to formulate quantum thermodynamics under the same paradigm underlying classical thermodynamics, which encompasses the coarse-graining nature of classical macroscopic measurements, a new framework has been recently proposed [20, 21]. The new theory proposes that when we limit our information about the system to that attainable by measurements of specific observables (thus not performing full state tomography), the full quantum state contains too much information, most of which is inaccessible to the observer. For example, if one is only able to perform energy measurements on the system, it is impossible to distinguish between any two states with the same energy or to obtain information regarding quantum coherences [20, 21].
To eliminate the redundancy created by this excess information, the theory introduces the concept of an emergent gauge symmetry. This symmetry is called emergent because the excess information contained in a quantum system is redundant only from the perspective of thermodynamics and not from information theory [20]. This is analogous to the classical case, where the coarse graining is implemented by ignoring the microscopic degrees of freedom that are not accessible to classical measurements. Building on the notion of redundant information introduced in Refs. [20, 21], density matrices, which act as carriers of information, are proposed to be interpreted as analogues of gauge potentials in gauge theories.
Although we can consider any set of observables, let us consider the case where we can only perform energy measurements on a quantum system described by a time-dependent density matrix $\rho_{t}$. This is the typical situation in many experimental platforms when the dimension of the system increases [22]. Then, for a general time-dependent Hamiltonian $H_{t}$, the mean energy of the system is
[…]
The main idea behind the theory is to identify a symmetry group such that any density operator linked to $\rho_{t}$ by the action of the group gives the same value for the mean energy. In this way, we have an equivalence class of systems such that
[…]
with $V_{t}$ representing the symmetry transformation, an element of the thermodynamic group. Since the matrices $V_{t}$ take one density matrix to another, we can assume they are unitary. Joining this fact with Eq. (2), we see that the set of transformations $V_{t}$ allowed must commute with the system’s Hamiltonian. This leads to a gauge group that is isomorphic to the following Lie group [21]:
$\mathrm{G}_{T}=\mathrm{U}(n_{t}^{1})\times\mathrm{U}(n_{t}^{2})\times...\times\mathrm{U}(n_{t}^{k}),$ (3)
where $\times$ denotes the Cartesian product and $\mathrm{U}(n_{t}^{i})$ is the unitary group of dimension $n_{t}^{i}$, the (in general, time-dependent) degeneracy of the $i$-th eigenvalue of the Hamiltonian. Thus, $\sum_{i}n_{t}^{i}=d$ is the dimension of the system Hilbert space. The group $\mathrm{G}_{T}$ is called the thermodynamic group. Note that the structure of the group will change if the set of observables is modified; however, the general construction remains the same. Gauge invariance is imposed here by stating that physical quantities, which are expressed as functions of the density matrix, must be invariant under the action of $\mathrm{G}_{T}$.
In further analogy with modern gauge theories, Ref. [20] introduces the notion of a covariant derivative $\mathrm{D}_{t}$ defined by a potential $A_{t}=i\dot{u}_{t}u_{t}^{\dagger}$, where $u_{t}$ is the unitary matrix that diagonalizes the chosen observable, the system’s Hamiltonian in this case. It is defined by the action
$\mathrm{D}_{t}(\cdot)=\partial_{t}(\cdot)+i[A_{t},(\cdot)],$ (4)
Based on this covariant derivative, one can write the gauge-invariant definitions of work $W_{inv}$ and heat $Q_{inv}$ for quantum systems [20] as
[…]
and
[…]
respectively. In these equations, the integration is performed over time, considering a thermodynamic process that occurs between times 0 and $\tau$. See Refs. [20, 21] for the physical meaning of these quantities. It is important to observe here that we have two distinct actions on the system. First, we have a physical process under which the state evolves, and we have the action of the thermodynamic group, which is just a symmetry action at each instant of time. Work and heat, as well as other physical quantities, are related to the physical process.
Another important thermodynamic functional that can be defined in a gauge-invariant fashion is the entropy, which takes the form [21]
$S_{\mathrm{G}_{T}}[\rho(t)]=S_{vn}[\rho^{E}(t)],$ (7)
[…]
In Eq. (8), the $\Pi_{n_{t}^{k}}$ are projectors onto the subspace associated with the degeneracy $n_{t}^{k}$, while $\mathbb{I}_{n_{t}^{k}}$ stands for the identity operator on that subspace. The index $E$ stands for the energy eigenbasis. If we choose a distinct observable, the structure of this operator will be the same, but the eigenprojectors will be those of the other observable. See Ref. [21] for further details.
Moreover, this procedure for constructing gauge-invariant quantities is not exclusive to entropy. Particularly, since $\rho^{E}$ is the average density matrix with respect to the thermodynamic group (via quantum twirling), it makes sense to define a general gauge-invariant thermodynamic functional as [21]
$F_{inv}[\rho(t)]=F[\rho^{E}(t)]$ (9)
as long as its non-gauge-invariant counterpart $F[\rho]$ is invariant under unitary transformations [21].
Thus, this theory defines physical quantities as those invariant under the action of the gauge group $\mathrm{G}_{T}$, in complete analogy to the gauge principle in the fundamental theories of physics. Here, we take a step forward in the formalisation of the theory by providing its constructions in terms of fibre bundles, the natural language in which modern gauge theories are formulated.
III Gauge theories as geometry of principal bundles
In modern mathematics and theoretical physics, fields are understood as sections of fibre bundles. This viewpoint makes explicit the geometric origin of local degrees of freedom, clarifies the role of gauge choices, and highlights how connections generalise ordinary directional derivatives. Fibre bundles, therefore, provide the natural mathematical framework for formulating such theories. For example, in gauge theories describing the fundamental interactions between elementary particles, the interaction is mediated by a gauge potential. Geometrically, a gauge potential is the local expression of a connection that lives on a principal fibre bundle over the four-dimensional spacetime.
The purpose of this section is twofold: first, to establish the notation used in the next section; and second, to give a concise overview of the main mathematical structures involved. Readers seeking a deeper treatment are referred to Refs. [7, 8], while Appendix A provides additional details.
Given two topological spaces $\mathds{E}$ (referred to as the total space) and $\mathds{B}$ (referred to as the base space), a Lie group $\mathrm{G}$ (the fibre), and a projection map $\pi\penalty 10000\ :\penalty 10000\ \mathds{E}\penalty 10000\ \rightarrow\penalty 10000\ \mathds{B}$, a principal $\mathrm{G}$-bundle $\xi=(\mathds{E},\pi,\mathrm{G},\mathds{B})$ is a geometric structure such that there exists an open covering $\{O_{\alpha}\}$ and a set of homeomorphisms $\phi_{\alpha}:\pi^{-1}(O_{\alpha})\rightarrow O_{\alpha}\times\mathrm{G}$ (called local trivializations of $\mathds{E}$) satisfying $\pi\phi_{\alpha}^{-1}(x,g)=x$, for $x\in O_{\alpha}$ and $g\in\mathrm{G}$ [10]. Figure 1 presents a schematic representation of these concepts.
Figure 1: A simple example of a bundle. The base space $\mathds{B}$ is the real line $\mathds{R}$, while the total space $\mathds{E}$ is the plane $\mathds{R}\times\mathds{R}$. The projection map $\pi$ associates a fiber $E_{p}$ to each point of $\mathds{B}$. The total space is the union of all the fibres $\mathds{E}=\cup_{p\in\mathds{E}}{E}_{p}$. The projection simply tells you which point of the base space a given fibre element belongs to. In physics, the base space is usually spacetime and the fibre encodes internal degrees of freedom attached to each spacetime point. The section $s$ of the bundle is also shown. A section is a choice of exactly one fibre element over each point of the base space. In physics, a section is the rule that assigns to each point in spacetime the value of a physical quantity living in the fibre.
The above definition reflects the idea that the total space of a principal bundle is locally, but not necessarily globally, a product space of $\mathds{B}$ and $\mathrm{G}$. In the case where $\mathds{E}$ is globally isomorphic to $\mathds{B}\times\mathrm{G}$, we say it is a trivial principal bundle [10]. We can also define a function $s\penalty 10000\ :\penalty 10000\ \mathds{B}\penalty 10000\ \rightarrow\penalty 10000\ \mathds{E}$ called a section of the bundle that locally inverts the projection $\pi$. See Fig. 1 for further details on this definition. A principal bundle is trivial if and only if there exists a global section; i.e., the projection can be globally inverted. The sections of principal fibre bundles in the context of gauge theories are associated with choices of gauge [10, 11].
In the geometric theory of gauge fields, one is generally interested in the following setup [7, 11, 8, 10]: First, in the principal $\mathrm{G}$-bundle $\xi$, the gauge group of the theory is $\mathrm{G}$ (e.g. $\mathrm{U}(1)$ for electromagnetism, $\mathrm{SU}(2)$ for the weak interaction, and $\mathrm{SU}(3)$ for the strong interaction). Most commonly, $\xi$ is the product bundle with $\mathds{B}=\mathcal{M}$, where $\mathcal{M}$ is $(3+1)$-dimensional Minkowski spacetime. This is because $\mathcal{M}$ is effectively $\mathbb{R}^{4}$ with a certain choice of metric; thus, it is contractible [10]. Since $\mathcal{M}$ is contractible, by Theorem 1 in Appendix A, it is always trivial; thus, $\xi$ is isomorphic to the product bundle, which is easier to work with in local coordinates.
[…]
Given $\xi$, we also have a vector bundle (a fibre bundle where the fibre has a vector space structure [10, 23]) called the associated bundle. The sections of this bundle are the particles of the theory [11].
The associated bundle of a principal bundle can be constructed in the following manner [23]: given $\xi$ and a linear representation $\rho$ of $\mathrm{G}$ acting on a vector space $\mathds{V}$, the associated bundle $\eta_{\rho}$ is the quotient of $\mathds{E}\times\mathds{V}$ by the following equivalence relation [23]:
$(p,\nu)\sim(p\cdot g,\rho(g^{-1})\nu),$ (10)
[…]
In a way, the entirety of classical electrodynamics is contained within the tensor $F$. The four Maxwell equations (in the absence of sources) can be elegantly written as the Yang-Mills equations
[…]
where $\star$ is the Hodge star operator. By analysing them component-wise, it can be directly seen that Eqs. (11) leads to all four Maxwell’s equations [7, 10, 8]. This sort of expression is advantageous since it is valid for an arbitrary base manifold and does not depend on any specific set of coordinates [8].
Lastly, it would be of interest to take a look at how charged particle fields arise in this formalism. Following the construction we did earlier for the associated bundle, we must choose a representation of the group $\mathrm{U}(1)$. All of the linear representations of the 1-dimensional unitary group are of the form $\rho(g)=e^{iqt}$, $q\in\mathbb{Z}$ [24]. Thus, choosing a representation is equivalent to choosing an integer. This integer $q$ is called the charge of the particle fields of the associated bundle. Individual matter fields of charge $q$ are then seen to be sections of this associated bundle [11].
The main goal of this work is to formalise the gauge theory of quantum thermodynamics [20, 21] in the language of fibre bundles, thus writing thermodynamics on the same mathematical basis as the fundamental theories of physics.
IV Geometric theory of quantum thermodynamics
We begin by addressing what might, at first, seem to be an inconsistency. The theory developed in Refs. [20, 21] presents a gauge theory of the thermodynamic group $\mathrm{G}_{T}$. On the other hand, the potential $i\dot{u}_{t}u^{\dagger}_{t}$, defined for the covariant derivative, takes values in the Lie algebra of the unitary group $\mathrm{U}(d)$, and not in that of the thermodynamic group. We will shortly see that this is not an inconsistency and that we are actually dealing with two distinct (although related) geometrical structures, and not just one.
Moreover, it is important to note that, since $\mathrm{G}_{T}$ is not a fundamental gauge group, such as the ones in the standard model, but instead emerges from a lack of information about the system at hand, we should not expect it to behave in the same way as the group $\mathrm{U}(1)$ does in electrodynamics, for example. Indeed, as will become clear later on, this emergent thermodynamic gauge has a rich geometric structure that sets it apart from those commonly studied in field theories.
To construct the relevant fibre bundles for the theory, we note that since all quantities of interest, like the density operator $\rho_{t}$ and the Hamiltonian $H_{t}$, depend only on time. Therefore, by analogy with the gauge theories of the standard model, the base space of our bundle should be time, which can be taken to be the real line $\mathds{R}$. Since $\mathds{R}$ is contractible [10], our bundle is necessarily trivial by Theorem 1 in Appendix A. As such, we take the liberty to employ the isomorphism between a trivial bundle and a product bundle of the base space and the fibre, and we always regard quantum thermodynamics as taking place on the product bundle.
For the fibre, seeing the definition of the covariant derivative, we take the entire unitary group $\mathrm{U}(d)$ acting on the Hilbert space of dimension $d$. We further note that, since $\mathrm{G}_{T}$ is a subgroup of $\mathrm{U}(d)$, any structure regarding the thermodynamic group can be regarded as a substructure of this principal $\mathrm{U}(d)$-bundle.
Thus, the first geometrical setup for quantum thermodynamics is a principal $\mathrm{U}(d)$ bundle $\xi=(\mathds{R}\times\mathrm{U}(d),\pi,\mathrm{U}(d),\mathds{R})$, with $\pi$ having the same meaning as before. We note that it makes sense for the gauge group to be $\mathrm{U}(d)$, as we are not currently concerned with any specific measurements of thermodynamic quantities that would require gauge invariance under the action of $\mathrm{G}_{T}$. Since this bundle arises from the unitary matrices $u_{t}$ that diagonalize the Hamiltonian $H_{t}$, we can observe that the gauge freedom associated with the unitary group is that of a choice of basis, which we use to work in the eigenbasis of the Hamiltonian.
The next step in characterising the bundle’s geometry is verifying that the potential present in the covariant derivative defines a connection form on $\xi$. This can be seen to be true by noting that if we slightly change the definition of the potential from $i\dot{u}_{t}u_{t}^{\dagger}$ to $A_{t}=\dot{u}_{t}u_{t}^{\dagger}$ (thus changing the covariant derivative to $\mathrm{D}_{t}\cdot=\partial_{t}+[A_{t},\cdot]$), we see that $A_{t}$ is precisely a right-invariant Maurer-Cartan form on $\mathrm{U}(d)$. Thus, by Theorem 2 in Appendix A, $A_{t}$ defines a connection on $\xi$.
Thus, we have thus far constructed a principal $\mathrm{U}(d)$-bundle $\xi$ and defined a connection on it. It is now time to understand how physical quantities such as $\rho_{t}$ and $H_{t}$ arise in this construction. To do this, we need to build an associated vector bundle.
We take $\mathds{V}$ to be the vector space of $d\times d$ Hermitian matrices and have the group $\mathrm{U}(d)$ act on it via the adjoint representation, such that $u\cdot\nu=u\nu u^{\dagger}$, where $u\in\mathrm{U}(d)$ and $\nu\in\mathds{V}$. Since our original principal bundle is trivial, we can always take the associated bundle to be the product vector bundle with total space $\mathds{R}\times\mathds{V}$. This means that each point on $\eta$ is a tuple $(t,\nu)$, where $t$ is an instance of time and $\nu$ is some $d\times d$ Hermitian matrix. By analogy with the notion of matter fields as sections of the associated bundle discussed earlier, and noting that a global section of $\eta$ associates each point $t\in\mathds{R}$ with a Hermitian matrix $\nu\in\mathds{V}$, we see that the "matter fields" of the gauge theory of quantum thermodynamics are simply time-dependent Hermitian observables.
Given the above argument, we naturally consider the density matrices not as emergent gauge potentials, as proposed in Refs. [20, 21], but as specific matter fields of the theory. Thus, in this framework, the thermodynamic quantities of interest arise as functionals of the matter fields (observables), much like in a quantum field theory several physical properties arise as functionals of matter fields (for example, Dirac or Klein-Gordon fields) [8, 9, 25].
With the construction we have so far, we can also see that, since the adjoint representation of a Lie group induces the adjoint representation of its Lie algebra, namely $A\cdot=[A,\cdot]$ [26], via differentiation, the covariant derivative (4) implements that precise action, thus keeping itself consistent with the rest of the structure.
We now return to the topic of the gauge symmetry associated with the thermodynamic group $\mathrm{G}_{T}$. As defined in Eq. (3), it is more accurately described as a time-dependent family of Lie groups. Thus, for different points in the base space $\mathds{R}$, $\mathrm{G}_{T}$ is not necessarily the same group. As such, we do not, in general, have a single principal $\mathrm{G}_{T}$-bundle that can be defined as a subbundle of $\xi$ with base space $\mathds{R}$. We can, however, note that at any instant $t$, we can form a trivial principal bundle with total space $\{t\}\times\mathrm{G}_{T}^{t}$, where by $\mathrm{G}_{T}^{t}$ we denote the thermodynamic group at time $t$. This bundle has a connection defined by the Maurer-Cartan form of $\mathrm{G}_{T}^{t}$, similarly to how we define the connection for the $\mathrm{U}(d)$ bundle. It is this structure that carries along the emergent gauge symmetry related to the redundant information contained in the density matrix. This can be seen by noting that our construction of the associated bundle applies identically in this case.
This second geometric structure defined above formalises the idea that the emergent gauge of quantum thermodynamics arises only concerning measurements, which we assume are taken at some specific time $t$. This can be understood as the geometric meaning of the coarse-graining introduced in Ref. [20]. In a situation where the degeneracies of the Hamiltonian (and consequently the thermodynamic group) do not change continuously in time but vary in specific times $\{t_{1},t_{2},...,t_{m}\}$, we can define, instead of bundles over points, bundles over time intervals $0,t_{1}),[t_{1},t_{2}),...[t_{m-1},t_{m}),[t_{m},\infty)$ equipped with connections given by the Maurer-Cartan forms and Theorem 2. This construction may be useful in a geometrical study of quench dynamics, where one turns on additional interaction terms in the Hamiltonian at a discrete set of times.
One important thing to note is that, since the curvature form is a 2-form and can also be written as the unique pullback of a 2-form in the base space [23], our principal bundles have globally vanishing curvature due to the fact that a 1-dimensional base space such as $\mathds{R}$ can have no non-vanishing 2-forms.
In an effort to make the discussion above easier to digest, we again evoke the example of electrodynamics for a final comparison, as shown in Table 1.
Electromagnetism Quantum Thermodynamics
$\mathrm{U}(1)$ symmetry $\mathrm{U}(d)$ and $\mathrm{G}_{T}$ symmetries
4-vector potential Maurer-Cartan connection
Electromagnetic field strength Vanishing curvature
Charged particle fields Hermitian observables
Table 1: Comparison between the geometrical structure of the gauge theories of electrodynamics and quantum thermodynamics.
V Discussion
In this work, we have mathematically formalized the gauge theory of quantum thermodynamics, whose main ideas were put forward in Refs. [20, 21]. Specifically, the theory was formulated through the geometry of a principal $\mathrm{U}(d)$-bundle with an additional geometric structure associated with measurements (namely, the thermodynamic group $\mathrm{G}_{T}$).
In our argument, we have separated which mathematical objects belong to each geometric structure associated with the theory. Furthermore, by constructing associated vector bundles, we were able to create an analogy between gauge-theoretic matter fields and time-dependent Hermitian observables in quantum thermodynamics. This result helps strengthen the link between the theory proposed in Refs. [20, 21] and the gauge theories in fundamental physics, such as Yang-Mills theory. In other words, we put thermodynamics in the same mathematical language as the fundamental theories in physics.
Furthermore, although the structure associated with the thermodynamic group does not, in general, form a single principal fibre bundle, it may form a collection of them if the degeneracies of the observables under consideration change discretely over time. In particular, formalising these ideas geometrically allows us to put forward a series of questions regarding how the topological properties of the thermodynamic group influence the thermodynamics of a quantum system.
One natural example of a question that arises from this theory is whether the homology and cohomology of the thermodynamic group can lead to any insight into the possible dynamics of a system. Since the thermodynamic group is a product of Lie groups, it should be possible to study this issue by employing Kunneth’s theorem [10].
Another perhaps less obvious direction is the study of relative homology and cohomology [10]. Since the thermodynamic group is a Lie subgroup of the unitary group $\mathrm{U}(d)$, one may construct relative homologies and cohomologies between them. This may be of interest, as both $\mathrm{G}_{T}$ and $\mathrm{U}(d)$ are relevant gauge groups of the theory. In particular, since relative Rham cohomology relates to differential forms on $\mathrm{U}(d)$ that vanish on $\mathrm{G}_{T}$, and gauge-invariant quantities in this framework carry integrals over the thermodynamic group, it makes sense that it may have some meaning in the theory.
Additionally, we know that in the theory of classical thermodynamics, there are many important inequalities, such as inequalities constraining changes in the free energies of a system [1]. Meanwhile, in topology, there are known relations between the critical points of functions and the topology of their domain. Namely, Morse theory [10] does so by constructing inequalities that relate the numbers of critical points with different indices of the function to the Betti numbers (ranks of homology and cohomology groups) of the space. From this perspective, we believe that Morse theory may offer insight into quantum thermodynamic equilibrium by examining the critical points of certain properly defined free energies. We also believe that it may be possible for a generalised Clausius inequality, such as the one hypothesised in [21], to have some relation to the structure of the thermodynamic group, whether through its algebraic or geometric/topological properties.
One last, perhaps less speculative direction to follow relates to the study of the thermodynamic group in the context of Lie group bouquets [27]. As discussed earlier, due to its time dependence, the thermodynamic group by itself doesn’t form a proper fibre bundle for all times (and thus doesn’t define a gauge theory). The concept of a Lie Group Bouquet, introduced in the seminal 1966 paper by Douady and Lazard [27], formalises the notion of a family of Lie groups parametrised by some other manifold (which, in this case, would be $\mathds{R}$). It is immediately obvious that the thermodynamic group forms one such bouquet, and it should be worthwhile to study it as such.
VI Acknowledgments
LCC acknowledges CNPq through grant 308065/2022-0, the financial support of the National Institute of Science and Technology for Applied Quantum Computing through CNPq grant 408884/2024-0, and the warm hospitality of the International Institute of Physics, where this work was developed with the support of the Simons Foundation (Grant No. 1023171, R.C.).
Appendix A Mathematical background
We briefly review the notion of Lie groups and fibre bundles. We focus on the topics that we employed in the main part of the text. The goal is to make the article more self-contained and accessible to researchers interested in quantum thermodynamics, information theory, and gauge theories.
The theory of Lie groups is discussed in Refs. [24, 23], while nice references on fibre bundles are [7, 28, 10]. See also [11, 8] for some topics related to principal bundles. We direct the interested reader to these references for a deeper treatment.
A.1 Bundles and fibre bundles
We provide a brief description of the main mathematical structure that will be employed to build a geometric theory of quantum thermodynamics.
A bundle is a mathematical structure that generalises the notion of a product space. More specifically, one can understand a bundle as a triple $(\mathds{E},\pi,\mathds{B})$, with $\mathds{E}$ being a topological space called the total space, $\mathds{B}$ a topological space called the base space, and $\pi:\mathds{E}\rightarrow\mathds{B}$ the projection map.
For every point $b\in\mathds{B}$, we call the space $\pi^{-1}(b)$ the fibre of the bundle over $b$. It is often useful to think of a bundle as a set of spaces (fibres) parametrised by the base space and glued together by the topology of the total space.
Bundles are the basic mathematical objects over which many other, more complex structures, such as vector bundles, fibre bundles, and principal bundles, are built. In this work, we are mainly interested in principal bundles. It is, however, often pedagogical to build the notion of a principal bundle from that of a fibre bundle. Thus, we define a fibre bundle by adding two elements to the bundle. A topological space $\mathds{F}$, called the fibre, and a group $\mathrm{G}$ of homeomorphisms of the fibre $\mathds{F}$, called the structure group. Moreover, there is an open covering $\{O_{\alpha}\}$ of $\mathds{B}$ such that, at each $O_{\alpha}$, $\pi^{-1}(O_{\alpha})$ is homeomorphic to the Cartesian product $O_{\alpha}\times\mathds{F}$. The homeomorphisms $\phi_{\alpha}:\pi^{-1}(O_{\alpha})\rightarrow O_{\alpha}\times\mathds{F}$ are called local trivializations of $\mathds{E}$.
Thus, one can see a fibre bundle as a topological space that is locally (but generally not globally) a product space of $\mathds{B}$ and $\mathds{F}$. If a fibre bundle $\xi$ is globally a product space (i.e, if $\mathds{E}=\mathds{B}\times\mathds{F}$), then $\xi=(\mathds{B}\times\mathds{F},\pi,\mathds{F},\mathrm{G},\mathds{B})$ is called a product bundle. A fibre bundle $\zeta=(\mathds{E}^{\prime},\pi^{\prime},\mathds{F},\mathrm{G},\mathds{B})$ is called trivial if it is isomorphic to the product bundle $\xi$.
A section of a fibre bundle is a map $s:\mathds{B}\rightarrow\mathds{E}$ that inverts the projection onto the base space. More specifically, we have $p\circ s=\mathds{1}_{\mathds{B}}$, where $\mathds{1}_{\mathds{B}}$ is the identity map on $\mathds{B}$. The condition that a bundle $\xi$ is trivial is equivalent to the existence of a global section on $\xi$.
We now state an important theorem:
Theorem 1.
Every fibre bundle over a contractible base space $\mathds{B}$ is trivial.
By contractible, we mean that the identity map on $\mathds{B}$ is homotopic to some constant map. In other words, the topological space can be continuously deformed into a single point.
We do not prove this theorem here since it would require many concepts that are outside the scope of this work. For those interested in the proof, Ref. [10] provides an intuitive outline of why Theorem 1 is true, while Ref. [29] provides a rigorous proof using homotopy theory.
A crucial example of a (generally non-trivial) fibre bundle associated with a manifold $\mathds{M}$ is the tangent bundle $\mathrm{T}\mathds{M}$, defined as
$\mathrm{T}\mathds{M}=\bigcup\limits_{p\in\mathds{M}}\mathrm{T}_{p}\mathds{M},$
where $\mathrm{T}_{p}\mathds{M}$ is the tangent space of $\mathds{M}$ at the point $p\in\mathds{M}$. A manifold $\mathrm{M}$ is called parallelizable if its tangent bundle is trivial [24].
A.2 Lie Theory
A Lie group $\mathrm{G}$ is a group that is also a differentiable manifold in such a way that the map
$p:(g,h)\in\mathrm{G}\times\mathrm{G}\rightarrow gh\in\mathrm{G}$ (12)
is smooth. In particular, as a consequence of the condition above, the operation of taking an inverse in a Lie group is also differentiable (more precisely, it is a diffeomorphism). Since both the operations of taking a product and an inverse are differentiable, they are continuous; thus, all Lie groups are topological groups.
The Lie algebra $\mathfrak{g}$ associated with a Lie group $\mathrm{G}$ is the space of vector fields invariant under the action of $\mathrm{G}$, with a bracket given by the Lie bracket of vector fields. For our purposes, however, it is more convenient to exploit the fact that the Lie algebra of a Lie group is isomorphic to $\mathrm{T}_{e}\mathrm{G}$, the tangent space of the group at the identity element $e$, and to define the Lie algebra directly as such. Although the Lie algebra is isomorphic to the tangent space at the identity, there is a more general relation between Lie algebras and the tangent spaces of their associated Lie groups, namely:
$\mathrm{T}_{g}\mathrm{G}\cong\mathrm{G}\times\mathfrak{g}.$ (13)
Since $\mathfrak{g}\cong\mathrm{T}_{e}\mathrm{G}$, this also means that every Lie group is parallelizable.
The triviality of the tangent bundle of a Lie group $\mathrm{G}$ allows us to define the so-called left- and right-invariant Maurer-Cartan forms:
[…]
and
[…]
and
[…]
Maurer-Cartan forms play an important role in the gauge theory of the thermodynamic group presented in the main part of the paper.
Given a Lie group $\mathrm{G}$ and a manifold $\mathrm{M}$, a smooth right action of $\mathrm{G}$ on $\mathrm{M}$ is a smooth map $\mu\,:\,\mathds{M}\times\mathrm{G}\rightarrow\mathds{M}$, which we will denote $x\cdot g=\mu(x,g)$, such that:
  1. 1.
    $x\cdot e=x$;
  2. 2.
    $(x\cdot g)\cdot h=x\cdot(gh)$
$\forall x\in\mathds{M}$, $g,h\in\mathrm{G}$. $e$ is the identity element in $\mathrm{G}$. We say that an action is free if $x\cdot g=x\implies g=e$.
Given two manifolds $\mathds{M}$ and $\mathds{N}$ where $\mathrm{G}$ acts on the right, a map $f\,:\,\mathds{M}\rightarrow\mathds{N}$ is said to be $\mathrm{G}$-equivariant if $f(x\cdot g)=f(x)\cdot g$, where $(x,g)\in\mathds{M}\times\mathrm{G}$.
A.3 Principal Bundles
Principal bundles are among the most important bundles that arise in physics, as they provide the mathematical structure on which one can formalise gauge theories as geometrical constructs.
A principal $\mathrm{G}$-bundle is a fibre bundle for which the structure group (a Lie group $\mathrm{G}$) and the fibre are one and the same. More specifically, given a Lie group $\mathrm{G}$, we define a principal $\mathrm{G}$-bundle $(\mathds{E},\pi,\mathrm{G},\mathds{B})$ as a smooth fibre bundle $(\mathds{E},\pi,\mathds{F},\mathrm{G},\mathds{B})$ such that [23]:
  1. 1.
    $\mathds{F}=\mathrm{G}$;
  2. 2.
    $\mathrm{G}$ acts freely on the right on $\mathds{E}$;
  3. 3.
    The local trivializations $\phi_{\alpha}:\pi^{-1}(O_{\alpha})\rightarrow O_{\alpha}\times\mathrm{G}$ are $\mathrm{G}$-equivariant, where $\mathrm{G}$ acts on $O_{\alpha}\times\mathrm{G}$ by $(b,h)\cdot g=(b,hg)$.
The most important objects in the study of the geometry of a principal $\mathrm{G}$-bundle are connections. More specifically, given a principal $\mathrm{G}$-bundle $\xi=(\mathds{E},\pi,\mathrm{G},\mathds{B})$, we say that a $\mathfrak{g}$-valued $1$-form $\omega$ on $\mathds{E}$ is a connection if and only if
  1. 1.
    $\omega$ is $\mathrm{G}$-equivariant;
  2. 2.
    $\omega(A_{p}^{*})=A$, where
[…]
    for $A\in\mathfrak{g}$ and $p\in\mathds{E}$.
The $A_{p}^{*}$ in the above definition are vector fields on $\mathds{E}$ called the fundamental vector fields. Connections are very important in differential geometry because they allow us to define the notion of parallel transport in general spaces.
Given a connection $\omega$ on a principal $\mathrm{G}$-bundle, one can always construct a curvature form associated with it. That is, from $\omega$, one can make a $\mathfrak{g}$-valued 2-form that acts on vector fields $X$ and $Y$ over the total space $\mathds{E}$ in the following manner
[…]
where $\cdot$$\cdot$] is the Lie bracket of $\mathfrak{g}$.
The notion of a connection on a principal $\mathrm{G}$-bundle makes evident the significance of the Maurer-Cartan forms for the geometry of a Lie group. Namely, if we regard a Lie group $\mathrm{G}$ as the principal $\mathrm{G}$-bundle product with total space $\{x\}\times\mathrm{G}$, where $\{x\}$ is a single point, then $\theta_{g}^{R}$ is a connection on the Lie group regarded as a bundle.
We now turn our attention to a result on connections on trivial principal $\mathrm{G}$-bundles that can be found in [23, 30] and is useful in our construction.
Theorem 2.
Let $\xi=(\mathds{B}\times\mathrm{G},\pi_{1},\mathrm{G},\mathds{B})$ be a trivial principal $\mathrm{G}$-bundle, and $\theta_{g}^{R}$ be the right-invariant Maurer-Cartan form on $\mathrm{G}$. Define the projection $\pi_{2}:\mathds{B}\times\mathrm{G}\rightarrow\mathrm{G}$, $\pi_{2}(x,g)=g$. Then the pullback $\omega_{g}=\pi_{2}^{*}(\theta_{g}^{R})$ is a connection on $\xi$.
Proof.
First, we prove that $\omega_{g}$ is $\mathrm{G}$-equivariant. Let $(X_{x},X_{g})$ be a vector in the tangent space $\mathrm{T}_{(x,g)}\mathds{B}\times\mathrm{G}\cong\mathrm{T}_{x}\mathds{B}\oplus\mathrm{T}_{g}\mathrm{G}$ and $h\in\mathrm{G}$. Then $\omega_{g}(X_{x},X_{g})\cdot h=\pi_{2}^{*}\theta_{g}^{R}(X_{x},X_{g})\cdot h=\theta_{g}^{R}(\pi_{2}(X_{x},X_{g}))\cdot h=\theta_{g}^{R}(X_{g})\cdot h$. But since $\theta_{g}^{R}$ is a connection on $\mathrm{G}$ regarded as a bundle over a single point, $\theta_{g}^{R}(X_{g})\cdot h=\theta_{g}^{R}(X_{g}\cdot h)=\omega_{g}((X_{x},X_{g})\cdot h)$. Thus, $\omega_{g}$ is $\mathrm{G}$-equivariant.
Now let
[…]
be a fundamental vector field on $\mathds{B}\times\mathrm{G}$. Then
[…]
But this is exactly the Maurer-Cartan form applied to a fundamental vector field of $\mathrm{G}$. Thus
[…]
Since $\omega_{g}$ satisfies both axioms of a connection on a principal $\mathrm{G}$-bundle, it is a connection on $\xi$. ∎